# FALCON: Fairness Learning via Contrastive Attention Approach to Continual Semantic Scene Understanding in Open World

### v1

@@ 4.2 Ablation Study
Effectiveness of Loss Contributions Table 3 illustrates the contributions of proposed learning objectives. For the model without using visual grammar, we only use a single unknown cluster ($N_{U}=1$) and adopt the nearest cluster strategies to assign clusters of unknown pixels. By using only cross-entropy loss, the mIoU performance remains low due to catastrophic forgetting and background shift problems. Meanwhile, with our fairness clustering loss $\mathcal{L}^{\alpha}_{Cont}$, visual grammar model $\phi$, and the cluster regularizer $\mathcal{R}$, the mIoU performance has been significantly improved to $37.9\%$ and $36.4\%$ on ADE20K 100-50 and ADE20K 100-10, respectively. Moreover, our FALCON has significantly promoted the fairness of segmentation models illustrated by the mIoU improvement of major and minor groups.
Table 4: Effectiveness of Visual Grammar on ADE20K.
(a) ADE20K 100-50
0-100 101-150 all Major Minor
Nearest Cluster 44.3 11.5 33.4 51.5 24.3
Fixed $\phi$ 44.6 17.6 35.6 52.0 27.4
Adaptive $\phi$ 44.6 24.5 37.9 52.1 30.8
(b) ADE20K 100-10
0-100 101-150 all Major Minor
Nearest Cluster 40.1 14.3 31.5 48.7 22.9
Fixed $\phi$ 43.0 18.5 34.9 50.6 27.0
Adaptive $\phi$ 44.4 20.4 36.4 51.8 28.7
@@
MiB [3] 37.0 24.1 32.6 38.3 23.5 10.6 26.6 29.6 21.0 6.1 16.1 27.7
Mask2Former PLOP [19] 44.2 26.2 38.2 41.1 34.8 15.9 28.5 35.2 33.6 14.1 27.1 33.6
@@
Table 7: Comparisons on Cityscapes.
Effectiveness of Visual Grammar We evaluate FALCON under three settings, i.e., Nearest Cluster, Fixed $\phi$ pretrained ImageNet (without updating on each learning step), and Adaptive $\phi$ (with updating on each learning step). As in Table 4, the mIoU result using only the nearest cluster remains ineffective. Meanwhile, the adaptive visual grammar model updated at each learning step further boosts the mIoU performance and promotes fairness, i.e., increased by $4.5\%$ and $4.9\%$ on ADE20K 100-50 and ADE20K 100-10 compared to the nearest cluster approach. In addition, we study the impact of choosing the number of features $M$ in the visual grammar model in our supplementary. Fig. 6 illustrates the feature distributions of unknown classes (future class). As a result, our FALCON approach is able to model features of unknown classes into different clusters and produce better and more compact clusters compared to the one without Fairness Learning via Contrastive Attention.
Figure 7: Visualization of Our Results on ADE20K 100-50.
@@ 4.3 Comparison with Prior SOTA Methods
ADE20K Table 5 presents our experimental results using DeepLab-V3 and Transformer networks compared to prior CSS methods. Overall, our proposed approach has achieved the SOTA performance compared to prior methods. In particular, by using DeepLab-V3, our approach has achieved SOTA performance, i.e., the mIoU results of $37.9\%$ and $+36.4\%$ on ADE20K 100-50 and ADE20K 100-10 benchmarks, higher than prior FairCL [52]. Meanwhile, our approach using Transformer has outperformed the prior SOTA CoMFormer [4] model by $+3.5\%$, $+8.0\%$, and $+2.6\%$ on ADE20K 100-50, ADE20K 100-10, and ADE20K 100-5 respectively. In addition, our mIoU results on the initial classes remain competitive with the upper-bounded results because our method is able to well handle the fairness problem compared to the fully supervised learning approach. We also report our results on the ADE20K 50-50 benchmark in the supplementary. As in Fig. 7, FALCON produces better segmentation maps compared to prior methods.
@@
Pascal VOC Table 7 presents the results of our FALCON on Pascal VOC benchmarks. Our proposed approach has consistently achieved the SOTA performance on three benchmarks. In particular, compared to the prior FairCL [52] approach, our methods using DeepLab-V3 have improved the mIoU performance up to $73.50\%$, $69.83\%$, and $62.41\%$ on Pascal VOC 15-5, Pascal VOC 15-1, and Pascal VOC 10-1, respectively. Additionally, by using the better network backbone, i.e., Transformer, the performance of the segmentation model is also further improved. Our results have reduced the gap with the upper bound performance.
@@
Cityscapes Table 7 reports the performance of our approach using DeepLab-V3 compared to prior methods on three different settings of Cityscapes benchmarks, i.e., Cityscapes 11-5, Cityscapes 11-1, and Cityscapes 1-1. As shown in the experimental results, the performance of our methods has consistently outperformed prior FairCL [52] approach by $+3.78\%$, $+3.14\%$, and $+6.02\%$ on three benchmarks. Similar to our experiments on ADE20K and VOC, the better network brings higher results. These results have shown the effectiveness of FALCON on various benchmarks.
5 Conclusions
@@
This paper has presented a novel Fairness Learning via Contrastive Attention approach to CSS. In particular, the fairness contrastive clustering loss has been introduced to model both catastrophic forgetting and fairness problems. Then, the visual grammar model was presented to model the unknown classes. The experimental results on different benchmarks have shown the SOTA performance and fairness promotion of our proposed FALCON approach.
@@
Limitations Our study chose a set of learning hyper-parameters to support our theoretical analysis. However, it potentially consists of several limitations related to choosing learning parameters and cluster initialization. The details of our limitations are discussed in the supplementary. These limitations will motivate future research to further improve our Fairness Learning via Contrastive Attention.
Supplementary
1 Proof of Propositions 1 and 2
1.1 Proof of Proposition 1
Proposition 1: If the contrastive clustering loss $\mathcal{L}_{Cont}(;,\mathbf{c})$ achieve the optimal value, the enforcement $\ell_{i}$ between the feature and the cluster will converges to $\ell_{i}=L^{-1}$.
Proof: Let us consider the optimization of the Eqn. (4) in the paper as follows:
$\begin{split}&\min-\sum_{i=1}^{L}\log\frac{\exp(\mathbf{f}^{t}_{i}\times\mathbf{c})}{\sum_{\mathbf{f}^{\prime}}\exp(\mathbf{f}^{\prime}\times\mathbf{c})}=-\sum_{i=1}^{L}\log\ell_{i}\\ &\text{subject to}\quad\sum_{i=1}^{L}\ell_{i}=\ell\end{split}$ (9)
where $\ell$ is the total enforcement between features $\mathbf{f}^{t}_{i}$ and cluster $\mathbf{c}$. Then, the optimization of Eqn. (4) in the paper can be rewritten by using Lagrange multiplier as follows:
($\mathcal{L}\left(\{\ell_{i}\}_{i=1}^{L},\lambda\right)=-\sum_{i=1}^{L}\log\ell_{i}+\lambda(\sum_{i=1}^{L}\ell_{i}-\ell)$ (10)
where $\lambda$ is the Lagrange multiplier. Then, the contrastive clustering loss in Eqn. (4) in the paper achieves minimum if and only if:
($\begin{split}\frac{\partial\mathcal{L}\left(\{\ell_{i}\}_{i=1}^{L},\lambda\right)}{\partial\ell_{i}}&=-\ell_{i}^{-1}+\lambda=0\\ \frac{\partial\mathcal{L}\left(\{\ell_{i}\}_{i=1}^{L},\lambda\right)}{\partial\lambda}&=\sum_{i=1}^{L}\ell_{i}-\ell=0\\ \Rightarrow\mathcal{L}\left(\{\ell_{i}\}_{i=1}^{L},\lambda\right)&=-L\log\frac{\ell}{L}\end{split}$ (11)
As the total enforcement between features and the cluster is normalized, i.e., $\ell\in[0..1]$, the contrastive clustering loss ($\mathcal{L}\left(\{\ell_{i}\}_{i=1}^{L},\lambda\right)$ achieves minimum when $\log\ell=0\Rightarrow\ell=1$. Then, the enforcement between a single feature and the cluster will be equal to $\ell_{i}=\frac{\ell}{L}=L^{-1}$.
1.2 Proof of Proposition 2
Proposition 2: If the fairness contrastive clustering loss $\mathcal{L}^{\alpha}_{Cont}(;,\mathbf{c})$ achieve the optimal value, the enforcement $\ell_{i}$ between the feature and the cluster will converges to $\ell_{i}=(\alpha^{-1}+L)^{-1}$.
Proof: We first define the the enforcement between transitive vector $\mathbf{v}$ and the cluster $\mathbf{c}$ as $\ell_{\mathbf{v}}=\frac{\exp(\mathbf{v}\times\mathbf{c})}{\sum_{\mathbf{f}^{\prime}}\exp(\mathbf{f}^{\prime}\times\mathbf{c})}$. Then, let us consider the optimization of Eqn. (5) in the paper as follows:
$\begin{split}&\min-\sum_{i=1}^{L}\alpha\log\ell_{i}-\log\ell_{\mathbf{v}}\\ &\text{subject to}\quad\sum_{i=1}^{L}\ell_{i}+\ell_{\mathbf{v}}=\ell\end{split}$ (12)
Similar to Eqn. (9), Eqn. (12) can be reformulated via Lagrange multiplier as follows:
($\mathcal{L}\left(\{\ell_{i}\}_{i=1}^{L},\lambda\right)=-\sum_{i=1}^{L}\alpha\log\ell_{i}-\log\ell_{\mathbf{v}}+\lambda(\sum_{i=1}^{L}\ell_{i}+\ell_{\mathbf{v}}-\ell)$ (13)
Then, the fairness contrastive loss $\mathcal{L}^{\alpha}_{Cont}$ achieves minimum if and only if:
($\begin{split}\frac{\partial\mathcal{L}\left(\{\ell_{i}\}_{i=1}^{L},\lambda\right)}{\partial\ell_{i}}&=-\alpha\ell_{i}^{-1}+\lambda=0\\ \frac{\partial\mathcal{L}\left(\{\ell_{i}\}_{i=1}^{L},\lambda\right)}{\partial\ell_{\mathbf{v}}}&=-\ell_{\mathbf{v}}^{-1}+\lambda=0\\ \frac{\partial\mathcal{L}\left(\{\ell_{i}\}_{i=1}^{L},\lambda\right)}{\partial\lambda}&=\sum_{i=1}^{L}\ell_{i}+\ell_{\mathbf{v}}-\ell=0\\ \Rightarrow\mathcal{L}\left(\{\ell_{i}\}_{i=1}^{L},\lambda\right)&=-\alpha L\log\frac{\alpha\ell}{1+\alpha L}-\log\frac{\ell}{1+\alpha L}\end{split}$ (14)
As in Eqn. (14), the fairness contrastive learning loss ($\mathcal{L}\left(\{\ell_{i}\}_{i=1}^{L},\lambda\right)$ archives minimum when $\log\ell=0\rightarrow\ell=1$. Thus, the enforcement between the single feature the cluster will be re-balanced as $\ell_{i}=\frac{\alpha}{1+\alpha L}=(\alpha^{-1}+L)^{-1}$.
2 Implementation
Implementation Our framework is implemented in PyTorch and trained on four 40GB-VRAM NVIDIA A100 GPUs. The contrastive loss in our implementation is normalized with respect to the number of samples. These models are optimized by the SGD optimizer [2] with momentum 0.9, weight decay $10^{-4}$, and a batch size of $16$. The learning rate of the first learning step and the continual steps is set to $10^{-4}$ and $5\times 10^{-5}$ respectively. To update the cluster vectors $\mathbf{c}$, following prior work [52, 27, 25], we maintain a set of $500$ features for each cluster and update the clusters after $100$ steps with a momentum $\eta=0.99$. In our domain incremental experiments, all clusters are updated at each learning step by momentum update. The number of features selected for each cluster in the visual grammar model is set to $M=128$. The balanced weight of CSS objective $\lambda_{CL}$ and the cluster regularizer $\lambda_{C}$ is set to $1$. Following the common practices [52, 27], the margin between clusters $\nabla$ is set to 10.
Unknown Cluster Initialization As mentioned in the main paper, we adopt the DB-SCAN algorithm to initialize the clusters for unknown samples. In addition, to reduce the noise clusters and isolated clusters, we also merge several close clusters, i.e., if the distance between two clusters is less than the margin $2\nabla$, these will be merged into a single cluster where the new cluster center will be the means of these two merging cluster centers. By empirical observation, we have noticed that the number of unknown clusters initialized at each learning step, i.e., $N_{U}$ at the current learning step $t$, is not greater than 1.5$\times$ times of the remaining classes (i.e., $|\mathcal{C}^{t+1..T}|$) in the dataset, e.g., in our ADE20K 100-50 experiments, at the first learning step of $100$ classes, there are $68$ unknown clusters that have been initialized while there are $50$ remaining unknown classes in the dataset.
Cluster Assignment In our approach, we use our visual grammar model to assign the cluster for each feature representation. Theoretically, although there is a possibility that a feature could not be assigned to a cluster via the visual grammar model, we have empirically observed that this issue rarely happens in our approach. Indeed, since we initialize the known clusters via the DB-SCAN, it guarantees that for each feature, there is at least one cluster nearby that the feature representation should belong to. However, to preserve the integrity of our approach, for the outlier features in cases that cannot be assigned clusters via the visual grammar model, these outliers will be heuristically assigned to their closest clusters as similar to [27, 52].
Continual Learning Procedure Algorithm 1 illustrates the training procedure of our CSS approach.
Algorithm 1 CSS Procedure At Learning Step $t$
1: Learning Step $t$, Dataset $\mathcal{D}^{t}$, Visual Grammar $\phi(;\Theta_{t-1})$, and Segmentation Model $F(;\Theta_{t-1})$
2:Step 0: Extract features on $\mathcal{D}^{t}$ by $F(;,\theta_{t-1})$
3:Step 1: Initialize new known clusters for $\mathcal{C}^{t}$ of features extracted in Step 0
4:Step 2: Initialize potential unknown clusters of features extracted in Step 0
5:Step 3: Train CSS Model $F(;\theta_{t})$ on $\mathcal{D}^{t}$
6:Step 4: Extract features on $\mathcal{D}^{t}$ by $F(;,\theta_{t})$
7:Step 5: Train Visual Grammar Model $\phi(;,\Theta_{t})$ on current known clusters $\mathbf{c}$ and features extracted in Step 4
8:return $F(;\theta_{t})$ and $\phi(;\Theta_{t})$
3 Additional Experimental Results
3.1 Experiment Results of ADE20K 50-50 Benchmark
Table 8 presents the results of our method on the ADE20K 50-50 benchmark compared to prior methods. For fair comparisons, we use the DeepLab-V3 and Transformer in this experiment. As shown in the results, our proposed FALCON approach significantly outperforms prior methods. The results of our approach have reduced the gap with the upper bound result.
Table 8: Experimental results on ADE20K 50-50 Benchmark
ADE20K 50-50 (3 steps)
Network Method 0-50 50-150 all
MiB [3] 45.6 21.0 29.3
PLOP [19] 48.8 21.0 30.4
LGKD+PLOP [59] 49.4 29.4 36.0
DeepLab-V3 RCIL [60] 47.8 23.0 31.2
RCIL+LGKD [59] 49.1 27.2 34.4
FairCL [52] 49.7 26.8 34.6
FALCON 50.6 31.2 37.6
[…]
FairCL [52] 49.6 27.8 35.6
Transformer FALCON 53.0 36.8 42.2
[…]
3.2 Ablation Study
Effectiveness of Choosing Margin normal-$\nabla$ Table 9 studies the effectiveness of the value of margin $\nabla$ to the performance of our approach on ADE20K 100-50 and ADE20K 100-10 benchmarks. As shown in the results, the change of $\nabla$ also slightly influences the performance of the model. Since the margin defines the distance between two clusters, while the smaller value of the margin $\nabla$ could cause the incorrect cluster assignment of the features, the larger value of the margin $\nabla$ could produce the less compact clusters.
Table 9: Effectiveness of Choosing Margin $\nabla$
(a) ADE20K 100-50
0-100 101-150 all Major Minor
$\nabla=5$ 44.4 21.8 36.9 51.9 29.4
$\nabla=10$ 44.6 24.5 37.9 52.1 30.8
$\nabla=20$ 44.7 22.2 37.2 51.7 29.9
(b) ADE20K 100-10
0-100 101-150 all Major Minor
$\nabla=5$ 43.2 18.7 35.0 50.5 27.3
$\nabla=10$ 44.4 20.4 36.4 51.8 28.7
$\nabla=20$ 43.5 19.9 35.7 51.2 27.9
Effectiveness of Choosing Number of Features $M$ We study the impact of choosing the number of features $M$ in the visual grammar model. As in shown Table 10, the optimal performance of our approach is $M=128$. When the number of features selected is small ($M=96$), it does not have enough number of features to form the visual grammar so the model is hard to exploit the correlation among features and the cluster. Meanwhile, when we increase the number of selected features ($M=256$), the clusters will consist of many outlier features (the ones that do not belong to the cluster), thus being challenging for the visual grammar model to exploit the topological structures of the feature distribution.
Table 10: Effectiveness of Number of Features $M$ in a Cluster of Visual Grammar Model.
(a) ADE20K 100-50
0-100 101-150 all Major Minor
$M=96$ 43.0 19.6 35.2 50.5 27.5
$M=128$ 44.6 24.5 37.9 52.1 30.8
$M=256$ 43.6 21.6 36.3 51.0 28.9
(b) ADE20K 100-10
0-100 101-150 all Major Minor
$M=96$ 42.2 16.4 33.6 50.2 25.3
$M=128$ 44.4 20.4 36.4 51.8 28.7
$M=256$ 42.7 17.1 34.2 50.6 26.0
Effectiveness of Different Segmentation Networks To illustrate the flexibility of our proposed approach, we evaluate our proposed approach with different network backbones. Table 11 illustrates the results of our approach using DeepLab-V3 [7], SegFormer [56] with different backbones, i.e., ResNet-50, ResNet-101, MiT-B2, and MiT-B3. As shown in the performance, the more powerful the segmentation model is, the better performance of the model is. In particular, our approach has shown its flexibility since it consistently improves the performance of the segmentation model and achieves the SOTA performance on two different benchmarks, i.e., the performance of Transformer models achieves $41.9\%$, and $40.3\%$ on ADE20K 100-50, ADE20K 100-10, respectively.
Table 11: Effectiveness of Different Backbones on ADE20K.
(a) ADE20K 100-50
Backbone 0-100 101-150 all Major Minor
DeepLab-V3 R-50 44.3 15.2 34.7 51.5 26.4
R-101 44.6 24.5 37.9 52.1 30.8
[…]
MiT-B3 47.5 30.6 41.9 53.8 35.8
(b) ADE20K 100-10
Backbone 0-100 101-150 all Major Minor
DeepLab-V3 R-50 43.5 16.5 34.5 51.1 26.2
R-101 44.4 20.4 36.4 51.8 28.7
[…]
MiT-B3 47.3 26.2 40.3 54.0 33.4
4 Relation to Knowledge Distillation
Knowledge Distillation is a common approach to continual semantic segmentation [19, 4, 60, 5]. Prior work in clustering [52] has shown that the clustering loss is an upper bound of the knowledge distillation loss. Formally, the knowledge distillation loss can be formed as follows:
$\small\mathcal{L}_{distill}(\mathbf{x}^{t},F,\theta_{t},\theta_{t-1})=\mathcal{L}(\mathbf{F}^{t-1},\mathbf{F}^{t})$ (15)
where $\mathbf{F}^{t}$ and $\mathbf{F}^{t-1}$ are the feature representations extracted from the model at learning step $t$ and step $t-1$, respectively, and the metric $\mathcal{L}$ measure the knowledge gap between $\mathbf{F}^{t}$ and $\mathbf{F}^{t-1}$. Then, given a set of cluster $\mathbf{c}$, we consider the following triangle inequality of the metric $\mathcal{L}$ as follows:
$\small\begin{split}\forall\mathbf{c}:\quad\mathcal{L}(\mathbf{F}^{t},\mathbf{F}^{t-1})&\leq\mathcal{L}(\mathbf{F}^{t},\mathbf{c})+\mathcal{L}(\mathbf{c},\mathbf{F}^{t-1})\\ \Leftrightarrow\underbrace{\mathcal{L}(\mathbf{F}^{t},\mathbf{F}^{t-1})}_{\mathcal{L}_{distill}}&\leq\frac{1}{|\mathcal{C}^{1..T}|}\sum_{\mathbf{c}}\left[\underbrace{\mathcal{L}(\mathbf{F}^{t},\mathbf{c})}_{\mathcal{L}_{Cont}}+\mathcal{L}(\mathbf{c},\mathbf{F}^{t-1})\right]\end{split}$ (16)
At the computational time of Contrastive Clustering loss, the set of cluster vectors $\mathbf{c}$ is fixed (could be considered as constants). In addition, the features extracted at learning step $t-1$, i.e., $\mathbf{F}^{t-1}$, are constant due to the fix pre-trained model $\theta_{t-1}$. Therefore, without a strict argument, the distance $\mathcal{L}(\mathbf{c},\mathbf{F}^{t-1})$ could be considered as constant. Therefore, Eqn. (16) can be further derived as follows:
$\footnotesize\begin{split}\underbrace{\mathcal{L}(\mathbf{F}^{t},\mathbf{F}^{t-1})}_{\mathcal{L}_{distill}}&=\mathcal{O}\Bigg{(}\underbrace{\mathcal{L}\frac{1}{|\mathcal{C}^{1..T}|}}_{Constant}\sum_{\mathbf{c}}\Big{[}\underbrace{\mathcal{L}(\mathbf{F}^{t},\mathbf{c})}_{\mathcal{L}_{Cont}}+\underbrace{\mathcal{L}(\mathbf{c},\mathbf{F}^{t-1})}_{\text{Constant}}\Big{]}\Bigg{)}\\ &=\mathcal{O}\Bigg{(}\underbrace{\sum_{\mathbf{c}}\mathcal{L}(\mathbf{F}^{t},\mathbf{c})}_{\mathcal{L}_{Cont}}\Bigg{)}\\ \Rightarrow\mathcal{L}_{distill}(\mathbf{F}^{t-1},\mathbf{F}^{t})&=\mathcal{O}\left(\mathcal{L}_{Cont}(\mathbf{F}^{t},\mathbf{c})\right)\end{split}$ (17)
where $\mathcal{O}$ is the Big-O notation. Hence, from Eqn. (17), without lack of generality, we can observe that the Contrastive Clustering Loss is the upper bound of the Knowledge Distillation loss. Therefore, by minimizing the Contrastive Clustering Loss, the constraint of Knowledge Distillation is also maintained due to the property of the upper bound.
5 Discussion of Limitations
In our paper, we choose a specific set of hyper-parameters and learning approaches to support our hypothesis. However, our work could contain several limitations. First, choosing the scaling factor $\alpha$ could be considered as one of the potential limitations of our approach. In practice, when data keeps continuously growing, the pre-defined scaling factor $\alpha$ could not be good enough to control the fairness among classes. Our work focuses on investigating the effectiveness of our proposed losses to fairness, catastrophic forgetting, and background shift problems. Thus, the investigation of balance weights among losses has not been fully exploited, and we leave this experiment as our future work. Third, initializing the unknown clusters at each training step could potentially be room for improvement since the bad initial clusters could result in difficulty during training and updating these clusters and linking the unknown clusters learned in previous steps and new initial unknown clusters at the current learning steps have been yet fully exploited in our method. In addition, while our approach is designed for the DeepLab-V3 and Transformer segmentation networks [7, 56], the extensions of FALCON to mask-based segmentation networks [13, 12, 4] could be a potential next research for further performance improvement. These limitations could motivate new studies to further improve Fairness Learning via the Contrastive Attention Approach to continual learning in the future.

### v3

@@ 4.2 Ablation Study
Effectiveness af Forgetting Each Learning Step. Table 5 reports the results of our model evaluated at each learning step. The performance of our model remains stable at each learning step. In particular, the maximum performance drop at learning step $4$ of classes 111-120 due forgetting is only $0.8\%$. Meanwhile, at the last learning task, the mIoU results of old classes are well maintained, with a slight drop of 0.4% in classes 131-140. As shown in the result, our model has effectively maintained the performance to avoid catastrophic forgetting while enabling the capability of learning classes.
Table 5: Effectiveness of CSS Model At Each Learning Step.
ADE20K 100-10
0-100 101-110 111-120 121-130 131-140 141-150 all
Step 1 44.9 -- -- -- -- -- 44.9
Step 2 44.8 15.8 -- -- -- -- 42.2
Step 3 44.7 15.7 22.1 -- -- -- 40.3
Step 4 44.6 15.6 21.3 26.3 -- -- 39.2
Step 5 44.5 15.4 21.2 26.1 24.4 -- 37.9
Step 6 44.4 15.4 21.2 26.1 24.0 15.5 36.4
avg 40.1
@@
Mask2Former MiB [3] 37.0 24.1 32.6 38.3 23.5 10.6 26.6 29.6 21.0 6.1 16.1 27.7
PLOP [21] 44.2 26.2 38.2 41.1 34.8 15.9 28.5 35.2 33.6 14.1 27.1 33.6
@@
Figure 7: The Comparison of Qualitative Results on ADE20K Between Our Approach and RCIL [76].
@@ 4.3 Comparison with Prior SoTA Methods
ADE20K. Table 6 presents our experimental results compared to prior CSS methods. By using DeepLab-V3, our approach has achieved SoTA performance, i.e., the mIoU results of $37.9\%$ and $+36.4\%$ on ADE20K 100-50 and ADE20K 100-10 benchmarks, higher than prior FairCL [63]. Meanwhile, FALCON using Transformer has outperformed the prior CoMFormer [4] model by $+3.5\%$, $+8.0\%$, and $+2.6\%$ on ADE20K 100-50, ADE20K 100-10, and ADE20K 100-5. In addition, our mIoU results on the initial classes remain competitive with the upper-bounded results because FALCON can well handle the fairness problem compared to fully supervised learning. Although ECLIPSE [32] exhibits slightly better on ADE20K 100-5 by only $0.7\%$ (all) due to the better segmentation framework (i.e., Mask2Former), FALCON significantly outperforms ECLIPSE on ADE20K 100-50 and ADE20K 100-10. We also report our results of ADE20K 50-50 in the appendix. As in Fig. 7, FALCON produces better segmentation maps than prior methods.
@@
Pascal VOC. Table 8 presents our results on Pascal VOC benchmarks. FALCON has consistently achieved the SoTA performance on three benchmarks. In particular, compared to the prior FairCL [63] approach, our methods using DeepLab-V3 have improved the mIoU performance up to $73.50\%$, $69.83\%$, and $62.41\%$ on Pascal VOC 15-5, Pascal VOC 15-1, and Pascal VOC 10-1, respectively. Additionally, by using the better backbone, the performance is further improved, and the gap with the upper-bound result is reduced.
@@
Cityscapes. Table 8 reports the performance of our approach using DeepLab-V3 compared to prior methods on three different settings, i.e., Cityscapes 11-5, Cityscapes 11-1, and Cityscapes 1-1. As in Table 8, the performance of our methods has consistently outperformed prior FairCL [63] approach by $+3.78\%$, $+3.14\%$, and $+6.02\%$ on three benchmarks. Similar to our experiments on ADE20K and VOC, the better network brings higher results.
5 Conclusions and Limitations
@@
Conclusions. This paper has presented a novel Fairness Learning via Contrastive Attention approach to CSS. In particular, the fairness contrastive clustering loss has been introduced to model both catastrophic forgetting and fairness problems. Then, the visual grammar model was presented to model the unknown classes. The experimental results on different benchmarks have shown the SoTA performance and fairness promotion of our proposed FALCON approach.
@@
Limitations. Our study chose a set of learning hyper-parameters to support our theoretical analysis. However, it potentially consists of several limitations related to choosing learning parameters and cluster initialization. Further discussion of limitations is presented in the supplementary.
Acknowledgment This material is based upon work supported by the National Science Foundation under Award No. OIA-1946391. We also acknowledge the Arkansas High-Performance Computing Center for providing GPUs.